1.) y=-2/3x+4 by graphing 3 points
Answer by checkley71(8403)

SETTING Y=0 WE GET 0=-2/3X+4 OR -4=-2/3X OR -12=-2X OR X=6 THUS ONE POINT IS (6,0)
SETTING X=12 WE GET Y=-2/3(12)+4 OR Y=-8+4 OR Y=-4 THUS ANOTHER POINT IS
(12,-4)
SETTING Y=2 THEN 2=-2/3X+4 OR -2=-2/3X OR -6=-2X OR X=3 THUS ANOTHER POINT IS
(3,2)